# Artificial Human Intelligence: The role of Humans in the Development of Next Generation AI

### v1

@@ 4.1 Reverse engineering a human skill
One of the main pillars of brain-inspired research is to understand how hard it is to fully grasp the various emergent behaviors that leads to a form of intelligence in the brain. This question dates back to 1980s and the difficulty of adapting it depends on the developmental trajectory of that skill. The premise of the past work was that we typically expect the difficulty of reverse-engineering any human skill to be roughly proportional to the amount of time that skill has been evolving in living beings including animals [94]. For example one of the oldest human skills is perception which is largely unconscious and so appear to us to be effortless. Examples are abundant: Anything to do with attention, visualization, motor skills, tactile system, social skills and so on. The hypothesis is that we should expect skills that appear effortless to be difficult to reverse-engineer. The dual to this hypothesis is skills that require effort may not necessarily be difficult to engineer at all.
@@
Emerging skills developed later in the evolutionary trajectory such as mathematics, engineering, gaming, logic, and scientific reasoning pose challenges as they diverge from our evolutionary predispositions. These abilities have developed relatively recently in historical terms, undergoing refinement mainly through cultural evolution over the past few thousand years. Initially, many prominent researchers assumed that once they had grappled with the ”hard” problems, the seemingly ”easy” challenges of vision and commonsense reasoning would fall in place. However, the last decade in AI has proven it wrong. This misconception stemmed from the realization that these supposedly straightforward problems are, in fact, incredibly difficult, leading to the machines’ proficiency in solving problems like logic and algebra was irrelevant, as these tasks are exceptionally easy for them to handle [95, 16, 96].
4.2 What makes humans special and/or not so special?
@@
However, human cognitive system is extremely good at trying/searching (visual search) [98] and finding out things [^3] We will define “things” later in the context, predominantly referring to objects, entities, etc. (in other words, quite good at ontologies – a set of concepts and categories in a subject area or domain that shows their properties and the relations between them) leading to good performance in languages. Humans also have a sparse but adequate understanding of physical reality and causality which leads to data processing efficiencies in their generalization capacities. Due to this valuable feature, their learned world-models usually turn out to be reasonable and in congruent with the required tasks. They are also good at grounding the language with mental simulation. It is widely recognized that the way humans perceive the world is organized around objects - mostly up to 3rd dimension and some lausy understanding of the 4th dimension. These objects act as fundamental building blocks for a variety of higher-level cognitive processes, including language, planning, and reasoning. The notion of the world being comprised of separate parts that can be processed autonomously and combined in countless ways enables humans to make generalizations that go well beyond their direct experiences.
@@
In the context of mental simulation, the ladder of causation is pertinent to our discussion [99, 100]. We should consider our question with respect to the three rungs of the ladder: (1) The initial stage involves the extraction of meaningful correlations from real-world representations through a process of observation. This endeavor aims to estimate the conditional expectations of events, thereby facilitating the formation of crucial associations. (2) The subsequent stage entails engaging with the environment using an accurate model of the world which might be formed partially in the previous rung. Interactions in this phase result in alterations to the characteristics of objects that are inherently linked or correlated, either directly or through other dependent or confounding variables. Given that such interventions occur over time, they engender a sequence of actions and reactions conducive to causal reasoning. (3) The final stage of the ladder, termed counterfactual thinking, necessitates envisioning alternative scenarios, evaluating them, and adjusting accordingly. This aspect is distinguished by its requirement for imagining changes and their potential repercussions, without necessarily engaging in active intervention. Achieving counterfactuality underscores one of humanity’s most demanding cognitive abilities: predictive reasoning.
4.3 A case for Modularity and Robustness
@@
The brain network modularity can be defined as the degree to which functional brain networks (connectivity) are divided into special subnetworks. Modularity is assumed to be beneficial to develop a robust brain against internal as well as external perturbations. Biological individual differences seem to suggest the mandatory modular organizations for the robustness of persistent activity to various perturbations [101]. Modularity does not typically form in computational models with standard training regimens and architectures. However, modularity can be enforced through various approaches such as changing the training or the optimization process. Biomimetic training for instance is shown to demonstrate more robust behaviour under out-of-distribution scenarios. In addition, more biologically plausible optimizations provide robustness and seem to allow the emergence of invariant representations more quickly [102].
@@
4.4 Missing elements towards achieving Human-level AI
@@
5 Challenges and Perspectives in Human-Level AI Development
5.1 Challenges in the development of human-level AI
@@
One of the most fundamental issues in the deveopment of human-level AI is concerned with the identification of mechanisms that lead to commonsense reasoning in complex systems such as humans. For one thing, AI systems that possess a reasonable understanding of commonsense knowledge are expected to make sense of everyday situations and make intelligent decisions based on implicit knowledge that humans take for granted. As indicated in [27], common sense in AI are expected to build representations of core physical constructs such as time, space, causality, and establish interactions between physical objects and humans. Through the leverage of abstraction and compositionality, they should support reasoning techniques capable of handling complex and incomplete knowledge. More fundamentally, they also need to possess a human-inspired learning system (also see section 4.4) that continually incorporates new knowledge from various interactions and sources, enabling deep understanding and general intelligence. While carefully putting together bits and pieces, structural resemblences to human anatomy should also bring together explainability and interpretability components, which should naturally fall in place due to the joint consideration of learning and architecture. Creating AI that can explain its decisions and actions in a human-understandable manner is very important for building trust between machines and humans. The mutual trust is especially important in mission-critical applications such as healthcare and autonomous vehicles, where understanding the rationale behind AI-driven decisions can be a matter of life and death.
@@
AI systems must hold up ethical principles and societal values as part of their design and ensure that algorithms running on them adopt fair, transparent, and unbiased behavioral characteristics. This undertaking involves navigating issues such as privacy, accountability, and the long-term impact of AI on biological intelligence as well as employment and other dynamics of society at large. It is not a surprise that the next generation AI and its implementation will generate significant societal impacts, necessitating moderation and legislative regulations.
@@
Training foundation AI models and enhancing their ability to generalize knowledge across different domains and tasks through specialized fine tunings, also known as transfer learning, is another major challenge. This adaptation to novel circumstances involves enabling foundational AI systems to learn novel concepts from only a few samples of data (a.k.a. few-shot learning [124, 107]), leveraging previously acquired knowledge to solve new problems more efficiently. However, as is the case with modeling many cognitive functions of the brain with regards to real-world problem solving, there is no widely accepted mathematical model of transfer learning that connects out-of-distribution generalization to the minimum number of model parameters required for updates. Despite easy to articulate, accomplishing such a goal involves many genuine techniques that would make machines look like humans. Of such mechanisms, meta-learning [125] and self-supervised learning [126] are the two other critical areas of human-inspired research. Developing AI systems that can learn to learn, adapt, and improve over time through meta-learning techniques are desirable properties of the future generations of this technology. The overall goal of learning involves designing algorithms that can automatically discover effective strategies and apply them to new tasks, allowing AI to become more autonomous, capable and self-directed over time.
@@
Integrating AI systems with physical bodies (embodiment) and sensory capabilities to enable interaction with the physical world is another significant challenge towards human-level AI. This concept, known as embodied intelligence [127], involves understanding how intelligent behavior emerges from the complex interplay between an agent’s hardware components, implementation of the software logic and its dynamic environment. Achieving this integration is crucial for performing complex tasks requiring learning based on sensorimotor skills [128].
@@
5.2 Transferability of human-specific attributes to machines
In order to transfer a human specific attribute (Ethics, Aesthetics, Coinciousness, Morality, Humor, Causality or Trust), it is a priority to define these attributes using a precise language. Without a clear systematic description, it would be hard to comprehend and even have a common consensus on their meanings. However, creating a clear understanding entails improving the precision of their dictionary definitions [133], which might not be an easy task to perform in a mathematical sense. In fact, having precise definition for biological and metabolic events could sometimes be pretty hard to frame, which are analyzed historically using a branch of mathematics known as information theory [134]. In fact, imprecision in the development of human attributes could be beneficial in accurate and robust decision making processes.
5.2.1 The role of imprecise nature of perception in learning
@@
5.2.2 The effect of neuroscience studies for real-world stimuli
@@
Ultimately, this natural shift towards more realistic experimental paradigms holds the potential to transform our basic understanding of brain function, offering new insights into how neural circuits support adaptive behavior in the real world where the brain is constantly exposed to multi-modal sensory input. Such advances could pave the way for more effective interventions and technologies designed to support cognitive function in everyday life, from improving educational tools to developing brain-machine interfaces that operate seamlessly within the complexities of the real world [143, 144, 141].
5.3 Alignment of SoTA approaches with human-inspired/human-assisted AI
There is no doubt that today’s AI models are useful in closed and regulated application domains addressing successfully the specific task requirements. There are various angles of criticism directed at these systems, including algorithmic bias, ethical concerns, safety issues, overreliance on their use, ethics and their impact on creativity development, among others [145]. Rather than diving into the details of each critical point, which is what past research have heavily done so far, I will rather explore the alignment of current models in the context of hardware, architecture built on this hardware, and the learning schema that generates the software in the following subsections. We shall start considering the effect of the today’s hardware architectures (Von Neumann or parallel architectures of GPUs etc.) on the trajectory of AI system development.
5.3.1 The influence of the architecture of today’s computational models
@@
One key distinction between biological and digital intelligence is the strength of the coupling between the software implementing the aspects of intelligence and the available hardware this software is designed and built on. Abstraction between the layers in any complex system development allows modularity between system components, making design manageable but complicating debugging and failure tracing. Additionally, as the data volume increases and the models get chunkier, the implementation of digital intelligence seems to require significant resources, exponentially scaling the required power to support their learning. However, digital intelligence, with its parameter values, network links, and overall model descriptions, is highly efficient for knowledge storage and transfer across different digital hardware, reflecting its decoupled design nature. In otherwords, the learning process, despite being heavy and overwhelming, could be executed only once for all and then transferred/distributed for further learning through specializations, which also highlights the significance of foundation models. In contrast, human intelligence is tightly coupled with biological hardware, offering physiological advantages besides its efficient analog knowledge representation and processing. This tight coupling makes the overall learning process quite challenging, as each hardware needs to be trained from scratch using innate knowledge integrated into the architecture.
5.3.2 Model Architecture and Learning
@@
Biological and computational learning are mechanistically different. While the biological system strengthen or weaken neural connections based on experience, primarily through mechanisms like synaptic plasticity, computational systems simulate learning through scalar/vector weight values optimized by data, optimization and appropriate use of cost functions. Neither the data nor the optimization are necessarily bio-inspired. Among many other distinctions between biological and computational systems, optimization and locality of learning rates are two distinctive features of mechanisms employed in the learning process.
@@
Evolution as a computational approach to biological optimization refers to the application of evolutionary algorithms to solve complex and often constrained optimization problems. These algorithms mimic the mechanisms of mutation, crossover, and selection to iteratively improve solutions, leveraging biological principles to find optimal or near-optimal configurations in various domains, including genetics, neural networks, and ecological modeling. This approach capitalizes on the inherent efficiency and adaptability of evolutionary processes to address intricate challenges in biochemical mechanisms and computational biology. Although backpropagation is not necessarily required for networks to attain biologically interesting properties [159], it might be a good idea to leverage the main properties of ecosystems in building computational tools to solve complex problems [160, 159]. The possibility of inserting ecological concepts in the optimisation process is key in the development of novel bio-plausible hybrid systems.
One of the fundamental issues of computational intelligence is the non-local learning rules, which poses a challenge because it remains unclear how individual neurons and synapses in biological circuits access information about the global state of a circuit or a specific optimization target. The sources highlight the ongoing research focused on translating these global rules into local synaptic learning rules that rely solely on information locally available to neurons and synapses, such as pre- and post-synaptic spike timing [161]. Related to this regional learning, one of the stregths of biological networks is that they typically balance learning among multiple competing objectives, whereas computational schemes focus on optimizing for a single task. Future research could explore incorporating factors like energy consumption, wiring length, processing speed, learning efficiency, and robustness into ANN optimization to gain a more comprehensive understanding of biological circuit design.
5.3.3 Comments on building blocks: MLP, KAN and beyond
@@
The Multi-Layer Perceptron (MLP) [162] is a widely adopted model to simulate biological information processing through continuous approximation mechanisms. In other words, it is a crude representation of high-level functionalities of synapses and axons/dendrites in the human brain, making it easier to apply calculus rules and to implement learning on digital hardware with the available instruction sets. The generalized use cases of MLP consists of multiple layers of interconnected nodes, or so called neurons, where each neuron performs a weighted linear sum of its inputs followed by a non-linear activation function. Due to Universal Approximation Theorem (UAT) [163], this intuitive structure enables MLPs to approximate any arbitrary non-linear continuous function by learning and modeling complex patterns and relationships within data.
@@
Arnold-Kolmogorov (AK) theorem suggests that any multivariate function can be decomposed into a sum of compositions of univariate functions, offering a theoretical basis for designing networks known as Kolmogorov-Arnold Networks (KANs) [164]. These networks leverage the theorem to simplify the construction of neural networks capable of approximating complex functions with fewer parameters and potentially more efficient architectures in which weightening is replaced with individual spline-based activation functions.
@@
However, we need to realize that the UAT and the AK theorems are two alternatives for multivariate complex function approximations and have found distinct implementation styles, inspiring computational hardware designs that might emulate biological information processing better. Concordantly, MLPs and KANs represent two different but similar approaches to neural network architecture design as well. While MLPs focus on signal linearity, flexibility, and generalization capabilities with voluminous data, Kolmogorov-Arnold Networks aim at non-linearity at the edges, optimizing the representation and computational efficiency. Despite the potential of reducing the number of parameters, theoretically, we can make a few observations for these networks: (1) MLPs with learnable activation functions that could be formulated in terms of the sum of splines have the potential to approximate activations of KANs. (2) This theoretical equivalence leads to the conclusion that the main issues with MLP-based network design such as the curse of dimensionality are still applicable to the fundamentals of KANs.
@@
5.3.4 Texture/Shape Bias and Fine-Tuning in vision models:
In a relentless pursuit of building models that better align with human perception, researchers have been focusing on two key features: The bias and a subsequent fine-tunning for knowledge transfer. The past decade of research has clearly demonstrated that understanding the role of texture and shape bias in models like Vision Transformers (ViT) and ResNet, along with the impact of fine-tuning and augmentation techniques, is crucial to enhancing model accuracy and human-like error consistency.

### v2

@@ IV-A1 Reverse engineering a human skill
One of the main pillars of brain-inspired research is to understand how hard it is to fully grasp the various emergent behaviors that lead to a form of intelligence in the brain. This question dates back to the 1980s and the difficulty of adapting it depends on the developmental trajectory of that skill. The premise of the past work was that we typically expect the difficulty of reverse-engineering any human skill to be roughly proportional to the amount of time that skill has been evolving in living beings, including animals [94]. For example, one of the oldest human skills is perception, which is largely unconscious and so appears to us to be effortless. Examples are abundant: Anything to do with attention, visualization, motor skills, tactile system, social skills, and so on. The hypothesis is that we should expect skills that appear effortless to be difficult to reverse-engineer. The dual to this hypothesis is skills that require effort may not necessarily be difficult to engineer at all.
@@
Emerging skills developed later in the evolutionary trajectory, such as mathematics, engineering, gaming, logic, and scientific reasoning, pose challenges as they diverge from our evolutionary predispositions. These abilities have developed relatively recently in historical terms, undergoing refinement mainly through cultural evolution over the past few thousand years. Initially, many prominent researchers assumed that once they had grappled with the ”hard” problems, the seemingly ”easy” challenges of vision and commonsense reasoning would fall in place. However, the last decade in AI has proven it wrong. This misconception stemmed from the realization that these supposedly straightforward problems are, in fact, incredibly difficult, leading to the machines’ proficiency in solving problems like logic and algebra being irrelevant, as these tasks are exceptionally easy for them to handle [95, 17, 96].
IV-A2 What makes humans special and/or not so special?
@@
However, human cognitive system is extremely good at trying/searching (visual search) [98] and finding out things [^3] We will define “things” later in the context, predominantly referring to objects, entities, etc. (in other words, quite good at ontologies – a set of concepts and categories in a subject area or domain that shows their properties and the relations between them) leading to good performance in languages. Humans also have a sparse but adequate understanding of physical reality and causality which leads to data processing efficiencies in their generalization capacities. Due to this valuable feature, their learned world-models usually turn out to be reasonable and in congruent with the required tasks. They are also good at grounding the language with mental simulations. It is widely recognized that the way humans perceive the world is organized around objects - mostly up to 3rd dimension and some lausy understanding of the 4th dimension. These objects act as fundamental building blocks for a variety of higher-level cognitive processes, including language, planning, and reasoning. The notion of the world being composed of separate parts that can be processed autonomously and combined in countless ways enables humans to make generalizations that go well beyond their direct experiences.
@@
In the context of mental simulation, the ladder of causation is pertinent to our discussion [99, 100]. We should consider our question with respect to the three rungs of the ladder: (1) The initial stage involves the extraction of meaningful correlations from real-world representations through a process of observation. This endeavor aims to estimate the conditional expectations of events, thereby facilitating the formation of crucial associations. (2) The subsequent stage involves engaging with the environment using an accurate model of the world that might be formed partially in the previous rung. Interactions in this phase result in alterations to the characteristics of objects that are inherently linked or correlated, either directly or through other dependent or confounding variables. Given that such interventions occur over time, they engender a sequence of actions and reactions that are conducive to causal reasoning. (3) The final stage of the ladder, termed counterfactual thinking, necessitates envisioning alternative scenarios, evaluating them, and adjusting accordingly. This aspect is distinguished by its requirement for imagining changes and their potential repercussions, without necessarily engaging in active intervention. Achieving counterfactuality underscores one of humanity’s most demanding cognitive abilities: predictive reasoning.
IV-A3 A case for Modularity and Robustness
@@
The brain network modularity can be defined as the degree to which functional brain networks (connectivity) are divided into special subnetworks. Modularity is assumed to be beneficial to develop a robust brain against internal as well as external perturbations. Biological individual differences seem to suggest the mandatory modular organizations for the robustness of persistent activity to various perturbations [101]. Modularity does not typically form in computational models with standard training regimens and architectures. However, modularity can be enforced through various approaches such as changing the training or the optimization process. For instance, biomimetic training is shown to demonstrate more robust behaviour under out-of-distribution scenarios. In addition, more biologically plausible optimizations provide robustness and seem to allow the emergence of invariant representations more quickly [102].
@@
IV-A4 Missing elements towards achieving Human-level AI
@@
IV-B Challenges and Obstacles
@@
One of the most fundamental issues in the deveopment of human-level AI is concerned with the identification of mechanisms that lead to commonsense reasoning in complex systems such as humans. AI systems that possess a reasonable understanding of commonsense knowledge are expected to make sense of everyday situations and make intelligent decisions based on implicit knowledge that humans take for granted. As indicated in [27], common sense in AI are expected to build representations of core physical constructs such as time, space, causality, and establish interactions between physical objects and humans. Through the leverage of abstraction and compositionality, they should support reasoning techniques capable of handling complex and incomplete knowledge. More fundamentally, they also need to possess a human-inspired learning system that continually incorporates new knowledge from various interactions and sources, enabling deep understanding and general intelligence. While carefully putting together bits and pieces, structural resemblences to human anatomy should also bring together explainability and interpretability components, which should naturally fall in place due to the joint consideration of learning and architecture. Creating AI that can explain its decisions and actions in a human-understandable manner is very important for building trust between machines and humans. The mutual trust is especially important in mission-critical applications such as healthcare and autonomous vehicles, where understanding the rationale behind AI-driven decisions can be a matter of life and death.
@@
AI systems must hold up ethical principles and societal values as part of their design and ensure that algorithms running on them adopt fair, transparent, and unbiased behavioral characteristics. This undertaking involves navigating issues such as privacy, accountability, and the long-term impact of AI on biological intelligence as well as employment and other dynamics of society at large. It is not a surprise that the next generation AI and its implementation will generate significant societal impacts, necessitating moderation, governance and legislative regulations.
@@
Training foundation AI models and enhancing their ability to generalize knowledge across different domains and tasks through specialized fine tunings, also known as transfer learning, is another major challenge. This adaptation to novel circumstances involves enabling foundational AI systems to learn novel concepts from only a few samples of data (a.k.a. few-shot learning [124, 107]), leveraging previously acquired knowledge to solve new problems more efficiently. However, as is the case with modeling many cognitive faculties of the brain with regards to real-world problem solving, there is no widely accepted mathematical model of transfer learning that connects out-of-distribution generalization to the minimum number of model parameters required for updates. Despite easy to articulate, accomplishing such a goal involves many genuine techniques that would make machines look like humans. Of such mechanisms, meta-learning [125] and self-supervised learning [126] are the two other critical areas of human-inspired AI research. Developing AI systems that can learn to learn, adapt, and improve over time through meta-learning techniques are desirable properties of the future generations of the AI technology. The overall goal of learning involves designing algorithms that can automatically discover effective strategies and apply them to new tasks, allowing AI to become more autonomous, capable and self-directed over time.
@@
Integrating AI systems with physical bodies (embodiment) and sensory capabilities to enable interaction with the physical world is another significant challenge towards achieving human-level AI. This concept, known as embodied intelligence [127], involves understanding how intelligent behavior emerges from the complex interplay between an agent’s hardware components, implementation of the software logic and its dynamic environment. Achieving this integration is crucial for performing complex tasks requiring learning based on sensorimotor skills [128].
@@
IV-C Transferability of human-specific attributes to machines
To transfer a uniquely human attribute—-such as ethics, aesthetics, consciousness, morality, humor, causality, or trust—it is essential to define these attributes using clear and specific language. Without a clear systematic description, it would be hard to comprehend and even have a common consensus on their meanings. However, creating a clear understanding entails improving the specificity of their dictionary definitions [133], which might not be an easy task to perform in a mathematical sense. In fact, having specific definition for biological and metabolic events could sometimes be pretty hard to frame, which are analyzed historically using a branch of mathematics known as information theory [134].
The difficulty of creating clear definitions is partly due to noisy nature of their evolutionary development. In fact, the imprecision in the development of such human attributes can be shown to be beneficial in accurate and robust decision making processes.
IV-C1 The role of imprecise nature of perception in learning
@@
IV-C2 The effect of neuroscience studies for real-world stimuli
@@
Ultimately, this natural shift towards more realistic experimental paradigms holds the potential to transform our basic understanding of brain function, offering new insights into how neural circuits support adaptive behavior in the real world where the brain is constantly exposed to multi-modal sensory input. Such advances could pave the way for more effective interventions and technologies designed to support cognitive function in everyday life, from improving educational tools to developing brain-machine interfaces that operate seamlessly within the complexities of the real world [143, 141].
Figure 7: Select AI Index technical perfrmance benchmarks v.s. human performance. Source: AI Index, 2024 (AI Index Report). Available from: https://aiindex.stanford.edu/report/
IV-D Alignment of SoTA approaches with HIns/HAss AI
There is no doubt that today’s AI models exceeded humans on a lot of intellectual task categories, as shown in Fig. 7. We notice that these models are useful in closed and regulated application domains addressing successfully the specific task requirements. There are various angles of criticism directed at these systems, including algorithmic bias, ethical concerns, safety issues, overreliance on their use, ethics and their impact on creativity development, among others [144]. Rather than diving into the details of each critical point, which is what past research have heavily done so far, I will rather explore the alignment of current models in the context of hardware, architecture built on this hardware, and the learning schema that generates the software in the following subsections. We shall start considering the effect of the today’s hardware architectures (Von Neumann or parallel architectures of GPUs etc.) on the trajectory of AI system development.
IV-D1 The influence of the architecture of today’s computational models
@@
One key distinction between biological and machine intelligence is the strength of the coupling between the software implementing the aspects of intelligence and the available hardware this software is designed and built on. Abstraction between the layers in any complex system development allows modularity between system components, making design manageable but complicating debugging and failure tracing. Additionally, as the data volume increases and the models get chunkier, the implementation of machine intelligence, the implementation of non-biological intelligence, seems to require significant resources, exponentially scaling the required power to support their learning. However, machine intelligence, with its parameter values, network links, and overall model descriptions, is highly efficient for knowledge storage and transfer across different digital hardware, reflecting its decoupled design nature. In otherwords, the learning process, despite being heavy and overwhelming, could be executed only once for all and then transferred/distributed for further learning through specializations, which also highlights the significance of foundation models. In contrast, human intelligence is tightly coupled with biological hardware, offering physiological advantages besides its efficient analog knowledge representation and processing. This tight coupling makes the overall learning process quite challenging, as each hardware needs to be trained from scratch using innate knowledge integrated into the architecture.
IV-D2 Model Architecture and Learning
@@
Evolution as a computational approach to biological optimization refers to the application of evolutionary algorithms to solve complex and often constrained optimization problems. These algorithms mimic the mechanisms of mutation, crossover, and selection to iteratively improve solutions, leveraging biological principles to find optimal or near-optimal configurations in various domains, including genetics, neural networks, and ecological modeling. This approach capitalizes on the inherent efficiency and adaptability of evolutionary processes to address intricate challenges in biochemical mechanisms and computational biology.
Although backpropagation is not necessarily required for networks to attain biologically interesting properties [158], it might be a good idea to leverage the main properties of ecosystems in building computational tools to solve complex problems [159]. The possibility of inserting ecological concepts in the optimisation process is key in the development of new bio-plausible hybrid systems.
Figure 8: Comparison between MLP v.s. KAN layer. While the edges of MLP just hold one weight, the edges of KAN hold more parameters because of splines.
IV-D3 Comments on building blocks (MLP, KAN and beyond)
@@
The Multi-Layer Perceptron (MLP) [160] is a widely adopted model to simulate biological information processing through continuous approximation mechanisms. In other words, it is a crude representation of high-level functionalities of synapses and axons/dendrites in the human brain, making it easier to apply calculus rules and to implement learning on digital hardware with the available instruction sets. The generalized use cases of MLP consists of multiple layers of interconnected nodes, or so called neurons, where each neuron performs a weighted linear sum of its inputs followed by a non-linear activation function as shown in Fig. 8. Due to Universal Approximation Theorem (UAT) [161], this intuitive structure enables MLPs to approximate any arbitrary non-linear continuous function by learning and modeling complex patterns and relationships within data.
@@
Arnold-Kolmogorov (AK) theorem suggests that any multivariate function can be decomposed into a sum of compositions of univariate functions, offering a theoretical basis for designing networks known as Kolmogorov-Arnold Networks (KANs) [162]. KAN networks leverage the AK theorem to simplify the construction of neural networks capable of approximating complex functions with fewer parameters and potentially more efficient architectures in which weightening is replaced with individual spline-based activation functions (see Fig. 8).
@@
UAT and AK theorems are two alternatives for multivariate complex function approximations and have found distinct implementation styles, inspiring computational hardware designs that might emulate biological information processing better. Concordantly, MLPs and KANs represent two different but similar approaches to neural network architecture design. While MLPs focus on signal linearity, flexibility, and generalization capabilities with voluminous data, Arnold-Kolmogorov Networks aim at non-linearity at the edges, optimizing the representation and computational efficiency. Despite the potential of reducing the number of parameters, theoretically, we can make a few observations for these networks: (1) MLPs with learnable activation functions that could be formulated in terms of the sum of splines have the potential to approximate activations of KANs. (2) This theoretical equivalence leads to the conclusion that the main issues with MLP-based network design such as the curse of dimensionality are still applicable to the fundamentals of KANs.
@@
IV-D4 Texture/Shape Bias and Fine-Tuning